Bwindi Impenetrable Forest (01.01.2010 to 03.31.2010)

You are here

PDF versionPDF version
January - March
Communications activities: 

ITFC/UWA annual information sharing workshop was attended on 23rd and 24th February 2010. All ITFC staff together with UWA senior staff attended this workshop. Tim O'Brien of WCS/TEAM also attended this workshop.The workshop was held at Lake Mburo National Park.

Tim O'Brien of WCS/TEAM visited ITFC and Bwindi. Tim visited one of the vegetation plots and the climate station site. Tim also advised us on several issues regarding TEAM activities implementation in Bwindi. ITFC directors (Douglas Sheil and Miriam van Heist) and Badru Mugerwa have held meetings with Uganda Wildlife Authority (UWA) officials. UWA officials include Mr. Charles Tumwesigye (Bwindi- Mgahinga Conservation Area (BMCA) chief manager) and Mr. Edwin Kagoda (warden Research and Monitoring BINP). Several issues were discussed including payments for UWA staff participating in TEAM activities. Badru Mugerwa has also held meetings with local leaders in areas where TEAM activities are being implemented. The meetings were geared towards smooth implementation of TEAM activities in Bwindi.

Unusual Events at the TEAM Site: 

Here in Bwindi, we have been receiving rainfall since December 2009. This period (December to February) is allocated as a dry season on the TEAM sampling sechedular.

New Species: 

No new species have been found and identified in BIF. Sampling for terrestrial vertebrates has not started in BIF. Plant species identification in vegetation plots is scheduled for April 2010.

Protocol Problems: 

We have experienced delays in the implementation of several activities. These include:

  • Regular rains we have been experiencing in Bwindi since December.
  • Delays in equipment shipment.
  • Long distances to reach some of the plots (made more difficult when coupled with terrain). We need two days to reach some of these plots.
Schedule Problems: 

Vegetation protocol All the six vegetation plots have been measured and tagged. A total of 3068 stems have been measured and tagged.Stems from three plots have been mapped. Species identification and specimen collection for all plots not yet done.We measured plots within four months (December - March) instead of the three months (December- February) allocated to this protocol on the online schedule. Climate monitoring protocol Climate data collection was scheduled to start in October. No monthly climate data is being collected. This is attributed to the delay in the shipment of equipment to the site. Climate station (tower and fencing) has been put in place. Razor wire was added on to the fencing. Some of the equipment has been received. Climate station will be assembled as soon as the rest of the equipment is received. Data collection will start immediately. Notes

  • We are still on schedule on the terrestrial vertebrate protocol. This activity is scheduled for May 2010.
  • We have made a kind request to the TEAM director of information systems to change our online schedule to reflect these changes.
Publications or Presentations in Scientific Meetings: